OpenStudy (anonymous):

A rectangle has a length of the fifth root of 16 inches and a width of 2 to the 1 over 5 power inches. Find the area of the rectangle. 2 to the 3 over 5 power inches squared 2 to the 4 over 5 power inches squared 2 inches squared 2 to the 2 over 5 power inches squared to me this looks like a foreign language!!!!! HELP?!

OpenStudy (anonymous):

\[\sqrt[5]{16}\times 2^{\frac{1}{5}}\] i guess is what they are asking for weird to mix up the radical and exponential notation

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

\[16^{\frac{1}{5}}\times 2^{\frac{1}{5}}\]

OpenStudy (anonymous):

since \(16=2^4\) this is \[\large 2^{\frac{4}{5}}\times 2^{\frac{1}{5}}\]

OpenStudy (anonymous):

???????????? im really sorry but math aint my strongest subject

OpenStudy (anonymous):

add up the exponents, get \[\huge 2^{\frac{4}{5}+\frac{1}{5}}=2^{\frac{5}{5}}=2^1=2\]

OpenStudy (anonymous):

so is it the second answer choice??

OpenStudy (anonymous):

or C

OpenStudy (anonymous):

it is C!!! 2 inches squared

OpenStudy (anonymous):

can you help me with a couple other ones?? Rewrite the radical as a rational exponent. the cube root of 2 to the seventh power

OpenStudy (anonymous):

\[\huge \sqrt[3]{2^7}=2^{\frac{7}{3}}\]

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

can you help with 3 more?? I don't want to bother but I only got an hour left to finish 2 tests

OpenStudy (anonymous):

\[\huge \sqrt[\color{red}n]{x^\color{blue}m}=x^{\frac{\color{blue}m}{\color{red}n}}\]

OpenStudy (anonymous):

k shoot

OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 5 to the fourth power, over 25 = 52 By simplifying 25 to 52 to make both powers base five, and subtracting the exponents By simplifying 25 to 52 to make both powers base five, and adding the exponents By finding the quotient of the bases to be, one fifth and cancelling common factors By finding the quotient of the bases to be, one fifthand simplifying the expression

OpenStudy (anonymous):

\[\frac{5^4}{25}=\frac{5^4}{5^2}=5^{4-2}=5^2=25\]

OpenStudy (anonymous):

i would go with By simplifying 25 to \(5^2\) to make both powers base five, and subtracting the exponents

OpenStudy (anonymous):

Rewrite the rational exponent as a radical expression. 3 to the 2 over 3 power, to the 1 over 6 power the sixth root of 3 the ninth root of 3 the eighteenth root of 3 the sixth root of 3 to the third power

OpenStudy (anonymous):

\[\huge (3^{\frac{2}{3}})^{\frac{1}{6}}\] like that ?

OpenStudy (anonymous):

i suppose it is then you get \[\large 3^{\frac{1}{9}}=\sqrt[9]{3}\]

OpenStudy (anonymous):

the sixth root of 3 the ninth root of 3 the eighteenth root of 3 the sixth root of 3 to the third power these are the answer choices

OpenStudy (anonymous):

go with the ninth root of three

OpenStudy (anonymous):

ok thnx Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 3 over 4 power, all over 2 to the 1 over 2 power the eighth root of 2 to the third power the square root of 2 to the 3 over 4 power the fourth root of 2 the square root of 2 that's the last one

OpenStudy (anonymous):

\[\huge \frac{2^{\frac{3}{4}}}{2^{\frac{1}{2}}}=2^{\frac{3}{4}-\frac{1}{2}}=2^{\frac{1}{4}}\]

OpenStudy (anonymous):

or \[\huge \sqrt[4]{2}\]

OpenStudy (anonymous):

omg!!!! thank you so much!!! I really owe you somehow!!!!!!!!

OpenStudy (anonymous):

lol good luck with the rest of your tests

OpenStudy (anonymous):

thanks and its 10/10 right

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

I gave you a whole bunch of medals

OpenStudy (anonymous):

thanks now if i could only redeem them for valuable prizes....

OpenStudy (anonymous):

yeah that would be great. you'd be rich then